A woman is 5 times as old as her youngest daughter. How old is each now if the sum of their ages is 50?

Answer by Alan3354(69443)

A woman is 5 times as old as her youngest daughter. How old is each now if the sum of their ages is 50?
---------------
d = daughter's age
d + 5d = 50
d = 50/6 = 25/3 yrs = 8 yrs 4 months
Mother's age = 125/3 yrs = 41 yrs 8 months